(1 point) A statistical investigation is conducted to see if the mean of a population of values is equal to 367, or Ho :,-367 HA : #367. A random sample of n = 136 values is taken from this population of values, producing a test statistic of Teale-一0.643 and a P-value of 0.5215. Choose the correct interpretation of the P-value. OA. If the null hypothesis is true, the chance of another sample producing stronger evidence against the null hypothesis is 0.5215. o B. If the null hypothesis is true, the probability of rejecting the null hypothesis is 0.5215. O C. From this sample, the null hypothesis is not to be rejected. D. From this sample, the null hypothesis is to be rejected.
Answer #1

P value =0.5215

If P value \leq \alpha ..................then we reject H0

If P value > \alpha .................then do not reject H0

Option C) is correct.